OpenStudy (anonymous):

consider the sequence 16,-8,4,-2,1, A. describe the pattern formed in the sequence. B. find the next three terms

OpenStudy (anonymous):

please help me please

terenzreignz (terenzreignz):

A geometric sequence with first term 16 and common ratio = 0.5

OpenStudy (lgbasallote):

do you see a pattern?

terenzreignz (terenzreignz):

I meant r = -0.5 My bad

OpenStudy (anonymous):

ugh no because i see it go from negative numbers to positive can you exsplain everything to me please?

terenzreignz (terenzreignz):

Ok... A geometric sequence is such a sequence that each term is the previous term multiplied by some constant known as the common ratio. In this case, can you work out the common ratio?

OpenStudy (anonymous):

okay let me see if i see one but yeah i know what a sequence is but like when it jumps from negative to positive im like lost

terenzreignz (terenzreignz):

Well, forget about that, for the moment... What's -8/16 ?

OpenStudy (lgbasallote):

\[\LARGE \text{common ratio} = \frac{\text{second term}}{\text{first term}}\]

terenzreignz (terenzreignz):

^I'd just like to add that that thing you said only applies if we're sure we're dealing with a geometric sequence :)

OpenStudy (anonymous):

-0.5

terenzreignz (terenzreignz):

That's right. Now what's 4/-8? what's -2/4? 1/-2?

OpenStudy (anonymous):

oh i get it now so thats the sequence right it you get your number if you divide that by the number right?

terenzreignz (terenzreignz):

You'll notice that each term is the previous term multiplied by -0.5... right? You do notice that, right? :)

OpenStudy (anonymous):

yeah okay now for the secound part

terenzreignz (terenzreignz):

Ok, with that established, what do you think comes after 1?

OpenStudy (anonymous):

okay so 1 / -0.5 right?

terenzreignz (terenzreignz):

No... 1 * -0.5 We established that each term is the previous term MULTIPLIED by -0.5, the common ratio. Try again :)

OpenStudy (anonymous):

oh crap my bad i was looking at my next question okay so 1 times -0.5 = -0.5 right

terenzreignz (terenzreignz):

That's right. What comes after -0.5, then? :)

terenzreignz (terenzreignz):

by the way, I think this problem wants fractions, not decimals, you might do well to put -(1/2) instead of -0.5

OpenStudy (anonymous):

so you still times it right? if so its -1

terenzreignz (terenzreignz):

No... what's -0.5 x -0.5? I'm pretty sure it's not -1

terenzreignz (terenzreignz):

try \[-\frac{1}{2}\times-\frac{1}{2}\]

OpenStudy (anonymous):

oh no i got it wrong hold on

OpenStudy (anonymous):

0.25?

terenzreignz (terenzreignz):

that's right 0.25, or 1/4 Now, finally, what comes after that?

OpenStudy (anonymous):

1/4 times -1/2 right?

terenzreignz (terenzreignz):

that's right :)

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

So what did you get?

OpenStudy (anonymous):

-0.345

terenzreignz (terenzreignz):

really? try it again, make sure

OpenStudy (anonymous):

0.125?

terenzreignz (terenzreignz):

Wait a sec, your 0.125, are you sure it's positive?

OpenStudy (anonymous):

no its negative

terenzreignz (terenzreignz):

watch your signs ... >.> Anyway, now you have your next three terms :)

OpenStudy (anonymous):

thax you :)

terenzreignz (terenzreignz):

To sum it all up, the sequence has the form : \[x_{n}=16\left( -\frac{1}{2} \right)^{n-1} = (-1)^{n+1}(16)(0.5)^{n-1}\]

terenzreignz (terenzreignz):

Where \[x_{n} \] is the nth term in the sequence